OpenStudy (anonymous):

Another confusing problem.. 7x + 2y = 10 -7x + y = -16

hartnn (hartnn):

Add the 2 equations, what u get ?

OpenStudy (anonymous):

Solving Systems by Elimination.

OpenStudy (anonymous):

Yeah..it's confusing. :/

hartnn (hartnn):

Add the 2 equations, can you ?

OpenStudy (anonymous):

just add downwards.

OpenStudy (anonymous):

x + 3y = -6? o.o

hartnn (hartnn):

ok, lets add 'x' terms 7x + (-7)x = (7-7)x =... ?

OpenStudy (anonymous):

0x..so x = 0?..

hartnn (hartnn):

no, its just 0x

OpenStudy (anonymous):

So... just "x".

OpenStudy (raden_zaikaria):